OpenStudy (anonymous):

Maybe I should have written it...6x (5+7)/4-7x2+3x3 I am getting 13, but my TEAS study manual says it is -72

OpenStudy (anonymous):

What's the problem?

OpenStudy (anonymous):

\[\frac{6 \times(5 + 7)}{4 - 7\times2 + 3\times3}\] This?

OpenStudy (anonymous):

6x(5+7)\div4-7x2+3x3

OpenStudy (anonymous):

If it's \[\frac{6 \times (5 + 7)}{4 - 7 \times 2 + 3 \times 3}\] the answer is -72 as your text book said but if it is \[\frac{6\times(5+7)}{4}-7\times2+3\times3\] the answer is 13 so what the actual answer is depends on the way it's written in the textbook

OpenStudy (anonymous):

Okay...it's written just as I wrote it...excep t where i have "div" they have a division symbol